# WaterSeeker: Pioneering Efficient Detection of Watermarked Segments in Large Documents

### v3

@@ Appendix B Pseudocode of WaterSeeker and Detection Baselines
1:procedure WinMaxDetection(tokens, $W_{min}$, $W_{max}$, threshold)
@@
6:              stat $\leftarrow$ CalculateStatistics(tokens[$i:i+W$])
@@
5:         stat $\leftarrow$ CalculateStatistics(tokens[$i:i+W$])
@@
Algorithm 3 WaterSeeker Algorithm
1:procedure SuspiciousRegionLocalization(tokens, $W$, $k$, $\theta_{1}$, $\theta_{2}$)
2:     Compute score list using sliding window
3:     Calculate $s_{\text{mean}}$, $s_{\text{top-k-mean}}$, and $s_{\text{std}}$
4:     Detect anomalies using threshold:
5:     $s_{\text{mean}}+\max((s_{\text{top-k-mean}}-s_{\text{mean}})\cdot\theta_{1},s_{\text{std}}\cdot\theta_{2})$
6:     Connect nearby outliers and filter short segments
7:     return filteredSegments
8:end procedure
9:procedure LocalTraverseDetection(tokens, suspiciousRegions, $W$, threshold)
10:     hasWatermark $\leftarrow$ False
11:     indices $\leftarrow$ [ ]
12:     for $(s^{\prime},e^{\prime})$ in suspiciousRegions do
13:         maxScore = $\infty$- ∞, maxIndice = ()
14:         for $s\in[s^{\prime},s^{\prime}+W)$ and $e\in(e^{\prime}-W,e^{\prime}]$ do
15:              if DetectWatermark(tokens[$s:e$])$>$ maxScore then
16:                  maxScore = DetectWatermark(tokens[$s:e$])
17:                  maxIndice = $(s,e)$
18:              end if
19:         end for
20:         if maxScore $>$ threshold then
21:              hasWatermark $\leftarrow$ True
22:              indices.append(maxIndice)
23:         end if
24:     end for
25:     return hasWatermark, indices
26:end procedure
27:procedure WaterSeeker(tokens, $W$, $k$, $\theta_{1}$, $\theta_{2}$, threshold)
28:     suspiciousRegions $\leftarrow$ SuspiciousRegionLocalization(tokens, $W$, $k$, $\theta_{1}$, $\theta_{2}$)
29:     hasWatermark, indices $\leftarrow$ LocalTraverseDetection(tokens, suspiciousRegions, W, threshold)
30:     return hasWatermark, indices
31:end procedure
Appendix C Detailed Proof for Aar
C.1 Expected p-value of watermarked text
In this sub-section, we aim to analyze how the expected p-value for watermarked text varies with changes in the detection window size $W$.
Recall the p-value calculation formula:
[…]
where $S=\sum_{i=1}^{W}\log(\frac{1}{1-u_{i}})$, and $W$ is the window size.
For watermarked tokens, $E[\log(\frac{1}{1-u_{i}})]=\mu_{1}$ , and for non-watermarked tokens, $E[\log(\frac{1}{1-u_{i}})]=\mu_{0}$, where $\mu_{1}>\mu_{0}$.
The expectation of $S$ for different W values is as follows:
  • When $W\leq L$: $E[S]=W\mu_{1}$
  • When $W>L$: $E[S]=L\mu_{1}+(W-L)\mu_{0}$
The expectation of p-value is:
$E[\text{p-value}]=E[\Gamma(S,W,\text{loc}=0,\text{scale}=1)]$ (13)
Since GammaTransform is non-linear, we cannot directly substitute $E[S]$. However, we can use Jensen’s inequality to obtain an approximation:
$E[\Gamma(S,W,0,1)]\geq\Gamma(E[S],W,0,1)$ (14)
Therefore, we can analyze $\Gamma(E[S],W,0,1)$ to obtain a lower bound for the expectation of p-value. Define function $f(W)=\Gamma(E[S],W,0,1)$:
  • When $W\leq L$: $f(W)=\Gamma(W\mu_{1},W,0,1)$
  • When $W>L$: $f(W)=\Gamma(L\mu_{1}+(W-L)\mu_{0},W,0,1)$
Analyzing the behavior of $f(W)$:
  • When $W\leq L$, both $E[S]$ and the shape parameter W increase as $W$ increases.
  • When $W$ just exceeds $L$, the growth rate of $E[S]$ suddenly decreases (from $\mu_{1}$ to $\mu_{0}$), while the shape parameter W continues to increase linearly.
Consider the behavior of $f(W)$ near $W=L$:
  • When $W$ increases from $L-\epsilon$ to $L$, $E[S]$ increases by $\epsilon\mu_{1}$, and the shape parameter increases by $\epsilon$.
  • When $W$ increases from $L$ to $L+\epsilon$, $E[S]$ increases by $\epsilon\mu_{0}$, and the shape parameter increases by $\epsilon$.
Since $\mu_{1}>\mu_{0}$, at the point $W=L$, the rate of decrease of $f(W)$ suddenly slows down. This suggests that $f(W)$ is likely to reach its minimum value at $W=L$, or at a point very close to $L$. While this analysis does not strictly prove that the expectation of p-value is minimized exactly at $W=L$, it strongly suggests that the expectation of p-value is likely to reach its minimum value at or very near $W=L$ [^1] For a more rigorous proof, a deeper mathematical analysis of the GammaTransform function or numerical simulations would be necessary.
C.2 Calculation of p-threshold to control false positive rate
Similar to the analysis for KGW method, we also need to analyze the constraints on the p-value threshold $p^{*}$ when the false positive rate within the specified window is set to be lower than a target value $\alpha$.
For non-watermarked text, $u_{i}\sim\text{Uniform}([0,1])$. Consequently, the test statistic $S$ follows a Gamma distribution: $S\sim\text{Gamma}(W,1)$, where $W$ is the shape parameter and 1 is the scale parameter. The p-value is calculated using the Gamma CDF:
[…]
where GammaCDF is the cumulative distribution function of the Gamma distribution with shape parameter $W$ and scale parameter 1. To achieve a false positive rate of $\alpha$, we need to set a threshold $p^{*}$ such that: $P(\text{p-value}<p^{*})=\alpha$.
Given the definition of p-value, this is equivalent to: $P(1-\text{GammaCDF}(S,W,1)<p^{*})=\alpha$, which can be rewritten as: $P(S>\text{GammaInv}(1-p^{*},W,1))=\alpha$, where GammaInv is the inverse of the Gamma CDF.
Since $S$ follows a $\text{Gamma}(W,1)$ distribution for non-watermarked text, we can express this as:
$1-\text{GammaCDF}(\text{GammaInv}(1-p^{*},W,1),W,1)=\alpha.$ (16)
@@ D.2 Setting the threshold to achieve a target false positive rate $\alpha$
For KGW, as analyzed in Section 5.1, when the window size is large, we can approximate using the Central Limit Theorem, resulting in $z^{*}=\Phi^{-1}(1-\alpha)$. When $\alpha=10^{-6}$, this gives $z\approx 4.75$. However, when the window size $W$ is small, the approximation to a normal distribution using the Central Limit Theorem may lead to significant deviations. Therefore, we will use the binomial distribution for precise calculations. $x\sim B(W,\gamma)$ describes the number of green tokens in a window of size $W$ follows a binomial distribution, therefore:
@@
For Aar, as analyzed in Appendix C, $p^{*}=\alpha$, therefore set $p^{*}=10^{-6}$ for all lengths.
@@ Appendix E Impact of Window Size on Watermark Intensity Calculation
We present a case study comparing watermark intensity calculations using window sizes $W=1,10,30,$ and $50$. The analysis encompasses the ground truth segment and 500 tokens on either side. Figures 4 and 5 illustrate the results for the KGW and Aar algorithms, respectively. The intensity curves reveal that small window sizes, particularly $W\leq 10$, introduce significant fluctuations. While $W=30$ exhibits reduced noise, it still presents instabilities, as shown in Figure 5(c) (the ground truth segment part). Overall, $W=50$ demonstrates the least noise. Consequently, we adopt $W=50$ for our main experiments.
Appendix F Sample Distribution in the Main Experiment
It can be observed from Figure 6 that all four datasets include samples of watermarked segments with varying intensities and lengths, demonstrating the comprehensiveness and fairness of the dataset construction.
(a) KGW, Llama
(b) Aar, Llama
(c) KGW, Mistral
(d) Aar, Mistral
Figure 6: Distribution of sample numbers for the four datasets involved in the main experiment.
Appendix G Further Experimental Results under Different Mixing Ratios
In main experiment, 100-400 watermarked tokens are inserted into non-watermarked text of 10,000 tokens, resulting in watermarked text mixing ratios of 1% to 4%. This section expands the evaluation by testing the performance of various methods across a broader range of mixing ratios, achieved by adjusting the length of the non-watermarked text. We deliberately avoid adjusting the mixing ratio by inserting multiple watermark segments into the non-watermarked text, as the WinMax method can only return the highest-scoring segment and is not designed for multi-segment insertion scenarios. Additionally, we do not use fixed-length watermark segments because in real-world detection scenarios, the length of the watermark segment is typically unknown. Moreover, if the watermark length were fixed and known, a simple Fixed-length Sliding Window detection method (with the window size equal to the watermark segment length) would suffice to solve the problem.
We provided further experimental results under three different settings:
  1. 1.
    100-400 watermarked tokens are inserted into 5,000 non-watermarked tokens, with a mixing ratio of 2% (100 / 5100) -7.4% (400 / 5400). The results are demonstrated in Table 6.
  2. 2.
    100-400 watermarked tokens are inserted into 2,000 non-watermarked tokens, with a mixing ratio of 4.8% (100 / 2100) -16.7% (400 / 2400). The results are demonstrated in Table 7.
  3. 3.
    100-400 watermarked tokens are inserted into 500 non-watermarked tokens, with a mixing ratio of 16.7% (100 / 600) -44.4% (400 / 900). The results are demonstrated in Table 8.
Table 6: Comparison of WaterSeeker with other baselines under the mixing ratio of 2%-7.4%.
Model Method KGW Aar
FPR $\downarrow$ FNR $\downarrow$ F1 $\uparrow$ IoU $\uparrow$ FPR $\downarrow$ FNR $\downarrow$ F1 $\uparrow$ IoU $\uparrow$
Llama-2-7b Full-text Detection 0.000 1.000 [0.990] 0.000 [0.020] 0.001 0.000 1.000 [0.957] 0.000 [0.083] 0.002
WinMax 0.003 0.317 0.810 0.621 0.010 0.380 0.761 0.565
FLSW-100 0.000 0.520 0.649 0.413 0.003 0.523 0.644 0.375
FLSW-200 0.000 0.457 0.704 0.404 0.000 0.560 0.611 0.350
FLSW-300 0.003 0.663 0.502 0.297 0.000 0.767 0.378 0.283
FLSW-400 0.003 0.867 0.235 0.230 0.000 0.920 0.148 0.205
WaterSeeker(Ours) 0.003 0.357 0.781 0.570 0.010 0.407 0.740 0.536
Mistral-7b Full-text Detection 0.000 1.000 [0.990] 0.000 [0.020] 0.000 0.000 1.000 [0.937] 0.000 [0.119] 0.004
WinMax 0.003 0.227 0.871 0.700 0.010 0.370 0.768 0.570
FLSW-100 0.000 0.503 0.664 0.443 0.000 0.543 0.627 0.364
FLSW-200 0.000 0.350 0.788 0.497 0.003 0.540 0.629 0.362
FLSW-300 0.000 0.527 0.643 0.392 0.003 0.763 0.382 0.269
FLSW-400 0.000 0.783 0.357 0.281 0.003 0.910 0.165 0.203
WaterSeeker(Ours) 0.000 0.243 0.861 0.676 0.010 0.383 0.758 0.548
Table 7: Comparison of WaterSeeker with other baselines under the mixing ratio of 4.8%-16.7%.
Model Method KGW Aar
FPR $\downarrow$ FNR $\downarrow$ F1 $\uparrow$ IoU $\uparrow$ FPR $\downarrow$ FNR $\downarrow$ F1 $\uparrow$ IoU $\uparrow$
Llama-2-7b Full-text Detection 0.000 1.000 [0.967] 0.000 [0.065] 0.005 0.000 1.000 [0.873] 0.000 [0.225] 0.018
WinMax 0.003 0.257 0.851 0.669 0.003 0.413 0.738 0.533
FLSW-100 0.000 0.473 0.690 0.440 0.003 0.570 0.600 0.350
FLSW-200 0.000 0.420 0.734 0.446 0.000 0.607 0.565 0.327
FLSW-300 0.003 0.630 0.539 0.339 0.000 0.777 0.365 0.257
FLSW-400 0.003 0.850 0.260 0.246 0.000 0.883 0.209 0.200
WaterSeeker(Ours) 0.003 0.287 0.831 0.642 0.003 0.443 0.714 0.501
Mistral-7b Full-text Detection 0.000 1.000 [0.980] 0.000 [0.039] 0.003 0.000 1.000 [0.870] 0.000 [0.230] 0.017
WinMax 0.000 0.220 0.876 0.710 0.003 0.353 0.783 0.590
FLSW-100 0.000 0.470 0.693 0.447 0.000 0.487 0.678 0.415
FLSW-200 0.000 0.360 0.780 0.497 0.003 0.500 0.665 0.371
FLSW-300 0.000 0.533 0.636 0.386 0.003 0.760 0.386 0.285
FLSW-400 0.000 0.747 0.404 0.288 0.003 0.887 0.203 0.212
WaterSeeker(Ours) 0.000 0.227 0.872 0.701 0.003 0.363 0.776 0.574
Table 8: Comparison of WaterSeeker with other baselines under the mixing ratio of 16.7%-44.4%.
Model Method KGW Aar
FPR $\downarrow$ FNR $\downarrow$ F1 $\uparrow$ IoU $\uparrow$ FPR $\downarrow$ FNR $\downarrow$ F1 $\uparrow$ IoU $\uparrow$
Llama-2-7b Full-text Detection 0.000 1.000 [0.787] 0.000 [0.352] 0.081 0.000 1.000 [0.693] 0.000 [0.469] 0.109
WinMax 0.000 0.247 0.859 0.685 0.000 0.373 0.770 0.568
FLSW-100 0.000 0.470 0.693 0.445 0.000 0.557 0.614 0.380
FLSW-200 0.000 0.383 0.763 0.461 0.000 0.553 0.618 0.340
FLSW-300 0.000 0.553 0.618 0.360 0.000 0.690 0.473 0.298
FLSW-400 0.000 0.740 0.413 0.265 0.000 0.837 0.281 0.231
WaterSeeker(Ours) 0.000 0.277 0.839 0.647 0.000 0.410 0.742 0.522
Mistral-7b Full-text Detection 0.000 1.000 [0.847] 0.000 [0.266] 0.061 0.000 1.000 [0.697] 0.000 [0.465] 0.112
WinMax 0.000 0.270 0.844 0.652 0.000 0.350 0.788 0.588
FLSW-100 0.000 0.510 0.658 0.432 0.000 0.530 0.639 0.393
FLSW-200 0.000 0.407 0.745 0.442 0.000 0.510 0.658 0.369
FLSW-300 0.000 0.570 0.601 0.352 0.000 0.690 0.473 0.294
FLSW-400 0.000 0.760 0.387 0.243 0.000 0.827 0.295 0.246
WaterSeeker(Ours) 0.000 0.303 0.821 0.612 0.000 0.407 0.745 0.543
Experimental results demonstrate that WaterSeeker exhibits high robustness across various mixing ratios, showing consistent performance trends across different settings. Specifically, WaterSeeker significantly outperforms full-text detection and FLSW methods in both detection accuracy and localization effectiveness, while achieving comparable performance to the upper bound method, WinMax.
Appendix H Further Application Scenario
Figure 7: An illustration of WaterSeeker applied in AI detection system.
WaterSeeker not only provides accurate detection results, but its localization capabilities also contribute to building a more transparent and interpretable AI detection system. As illustrated in Figure 7, the AI detection system powered by WaterSeeker receives documents for analysis and outputs three components: (1) Detection Result, which indicates whether AI assistance was utilized; (2) Suspicious AI-generated Segments, which highlights segments identified as potentially AI-generated; and (3) AI Ratio, which displays the proportion of content attributed to AI. For example, in the context of academic integrity, highlighting suspicious segments provides actionable evidence for assessments and appeals. Additionally, reporting the AI ratio allows for more flexible establishment of academic misconduct standards.

### v4

@@ Appendix C Pseudocode of Detection Baselines
1:procedure WinMaxDetection(tokens, interval, threshold)
@@
6:              stat $\leftarrow$ WatermarkScore(tokens[$i:i+W$])
@@
5:         stat $\leftarrow$ WatermarkScore(tokens[$i:i+W$])
@@ D.2 Setting the threshold to achieve a target false positive rate $\alpha$
KGW. For KGW, as analyzed in Section 5.1, when the window size is large, we can approximate using the Central Limit Theorem, resulting in $z^{*}=\Phi^{-1}(1-\alpha)$. When $\alpha=10^{-6}$, this gives $z\approx 4.75$. However, when the window size $W$ is small, the approximation to a normal distribution using the Central Limit Theorem may lead to significant deviations. Therefore, we will use the binomial distribution for precise calculations. $x\sim B(W,\gamma)$ describes the number of green tokens in a window of size $W$ follows a binomial distribution, therefore:
@@
Aar. For Aar, recall the p-value calculation formula:
[…]
where $S=\sum_{i=1}^{W}\log(\frac{1}{1-u_{i}})$, and $W$ is the window size. For non-watermarked text, $u_{i}\sim\text{Uniform}([0,1])$. Consequently, $S$ follows a Gamma distribution: $S\sim\text{Gamma}(W,1)$, where $W$ is the shape parameter and 1 is the scale parameter. Equation 17 is equivalent to:
[…]
where GammaCDF is the cumulative distribution function of the Gamma distribution with shape parameter $W$ and scale parameter 1. To achieve a false positive rate of $\alpha$, we need to set a threshold $p^{*}$ such that: $P(\text{p-value}<p^{*})=\alpha$. Given the definition of p-value, this is equivalent to: $P(1-\text{GammaCDF}(S,W,1)<p^{*})=\alpha$, which can be rewritten as: $P(S>\text{GammaInv}(1-p^{*},W,1))=\alpha$, where GammaInv is the inverse of the Gamma CDF. Since $S$ follows a $\text{Gamma}(W,1)$ distribution for non-watermarked text, we can express this as:
$1-\text{GammaCDF}(\text{GammaInv}(1-p^{*},W,1),W,1)=\alpha.$ (19)
@@ Appendix E Impact of Window Size on Watermark Intensity Calculation
We present a case study comparing watermark intensity calculations using window sizes $W=1,10,30,$ and $50$. The analysis encompasses the ground truth segment and 500 tokens on either side. Figures 5 and 6 illustrate the results for the KGW and Aar algorithms, respectively. The intensity curves reveal that small window sizes, particularly $W\leq 10$, introduce significant fluctuations. While $W=30$ exhibits reduced noise, it still presents instabilities, as shown in Figure 6(c) (the ground truth segment part). Overall, $W=50$ demonstrates the least noise. Consequently, we adopt $W=50$ for our main experiments.